hello I would like to ask for help setting up a word problem. I know how to solve it once it is set as an equation but I sometimes have trouble getting them in the right format to get the correct answer. Here is the problem:
A computer keyboard is 11 in. longer than it is wide. If the length and width are increased by 2 in. the area is increased by 54 in. Find the length and width of the keyboard.
Thank you very much I really appreciate the help. I like to figure these out by myself but can't this one

Answer by stanbon(75887)

A computer keyboard is 11 in. longer than it is wide. If the length and width are increased by 2 in. the area is increased by 54 in. Find the length and width of the keyboard.
-----
Initial dimensions:
width = x
length = x+11
Area = x(x+11) = x^2+11x
-------
Modified dimensions:
width = x+2
length = x+13
Area = (x+2)(x+13) = x^2 + 15x + 26
--------
Equation:
new area - old area = 54 sq in
x^2 + 15x + 26 - (x^2+11x) = 54
4x + 26 = 54
4x = 28
x = 7 inches (original width)
x+11 = 18 inches (original length)
=========
Cheers,
Stan H.
===================